# Numerical optimization for the compatibility constant of the lasso

### v1

@@ 3.2 Mixed-Integer Quadratic Programming
The problem (9) can be easily solved when the number of elements in active set, $s$, is sufficiently small. However, since the number of sign patterns to be (9) is $2^{s-1}$ and grows exponentially, the computation of (9) becomes infeasible when $s$ is large.
To avoid enumerating all $2^{s-1}$ sign patterns explicitly, we formulate the problem as a mixed-integer quadratic program (MIQP, BertsimasTsitsiklis1997; BertsimasWeismantel2005). For each $j\in S$, we introduce nonnegative variables $v_{j}^{+},v_{j}^{-}\geq 0$ and a binary variable $b_{j}\in\{0,1\}$ to express $v_{j}$ as follows:
$v_{j}=v_{j}^{+}-v_{j}^{-},\quad v_{j}^{+}\leq Mb_{j},\quad v_{j}^{-}\leq M(1-b_{j}),$ (10)
where $M>0$ is a sufficiently large constant. This formulation is referred to as Big-$M$ formulation (NemhauserWolsey1988). With this formulation, the constraint $\|\bm{v}_{S}\|_{1}=1$ can be written as a linear equality
@@
Alternatively, we may replace the Big-$M$ constraints $v_{j}^{+}\leq Mb_{j}$ and $v_{j}^{-}\leq M(1-b_{j})$ in (11) with the SOS1 (special ordered set of type 1) constraint on $(v_{j}^{+},v_{j}^{-})$, which enforces that at most one of them is nonzero (BealeTomlin1970; Williams2013).
@@ 4.1 Evaluation of compatibility constant via QP
For each iteration of the Monte Carlo simulation, we generate a dataset and compute the compatibility constant $\phi$ by solving the QP with the OSQP solver using the osqp package in R. The lasso estimator is implemented with glmnet. We obtain the MSE and its upper bound:
$\mathrm{MSE}=\frac{1}{n}\|\bm{X}(\hat{\bm{\beta}}-\bm{\beta})\|_{2}^{2},\qquad\mathrm{bound}=9\lambda^{2}s/\phi^{2}.$ (12)
@@
Figure 1 presents the compatibility constant $\phi$ for different $n$, $p$, and $\rho$ under $R=10$ runs. The results show that the compatibility constant becomes large as the number of observations increases. This behavior can be explained by the fact that the rank of the Gram matrix $\hat{\Sigma}=\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}/n$ is less than or equal to $\min(n,p)$. When $n$ is small, the Gram matrix $\hat{\Sigma}$ has low rank and a large null space, allowing vectors satisfying the condition $\mathcal{C}:=\{\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}\}$ to align with directions where the quadratic form $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ becomes nearly zero. As $n$ increases, the rank of $\hat{\Sigma}$ increases and its null space shrinks, and as a result, it becomes harder to find such directions within $\mathcal{C}$.
@@
We also observe that as $p$ increases, the compatibility constant decreases. This is probably because increasing $p$ enlarges the feasible set $\mathcal{C}$, which provides more degrees of freedom and makes it easier to find directions in which the quadratic form $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ is small. In particular, when $p$ is large and $n$ is small, the compatibility constant sometimes becomes $\phi=0$, which implies the compatibility condition does not hold. Therefore, for high-dimensional data, several hundreds of sample sizes would be necessary to get a reasonable compatibility constant. The compatibility constant also highly depends on the value of $\rho$; as the value of $\rho$ increases, the value of $\phi$ decreases. This can be explained by noting that the compatibility constant behaves similarly to the smallest eigenvalue of the covariance matrix, $1-\rho$, which decreases as $\rho$ increases.
@@
The red dashed line indicates the upper bound of compatibility constant for the population covariance matrix, $\Sigma$; when $s$ is even, $\phi^{2}=1-\rho$, and when $s$ is odd, $\phi^{2}\leq(1-\rho)\left(1+\frac{1}{s(p-s)}\right)$. For derivation of the above upper bound, please refer to Appendix A. We observe that when $p$ is large, the limit would converge to the true but slowly, so that we need a relatively large number of observations to obtain a reasonable bound derived from the compatibility constant of the population covariance matrix. In other words, the compatibility condition can fail even under moderate correlation structures, simply due to finite-sample limitations rather than the population correlation itself.
Figure 2: Scaled MSE and upper bound, ${\rm MSE}_{\rm scaled}$ and ${\rm bound}_{\rm scaled}$, for varying $(n,p,\rho)$.
Figure 2 shows the scaled mean squared error (MSE) and its upper bound as below
${\rm MSE}_{\rm scaled}=\frac{n}{\log p}{\rm MSE},\quad{\rm bound}_{\rm scaled}=\frac{n}{\log p}{\rm bound},$ (13)
@@
where the MSE and its bound are defined in (12). Here, the “scaled” means multiplication of $n/\log p$ so that the right-hand side of inequality (4), ${\rm MSE}\lesssim C\log p/n$, has a quantity of order one.
@@
The result shows that the scaled MSE is relatively stable regardless of the value of $n$ and $p$, except for small $n$. On the other hand, the upper bound is unstable when $n$ is small and $p$ is large. This is because the compatibility constant can be small for small $n$ as discussed in Figure 1. In particular, when $p$ is sufficiently large and $n$ is small, the upper bound cannot be obtained because of $\phi=0$. Therefore, the upper bound made by the small sample sizes can be much larger than that with large sample sizes.
@@
Figure 3: Ratio between the MSE and the upper bound, $\text{MSE}/\text{bound}$, for varying $(n,p,\rho)$.
@@
Figure 3 shows the ratio between the MSE and its upper bound. The ratio becomes extremely large for high-dimension and small sample sizes. In particular, the ratio can be $10^{5}$ to $10^{6}$, which implies the upper bound does not make sense to evaluate the MSE. Therefore, the compatibility constant cannot be useful unless $n$ is relatively large. For large sample sizes, the ratio is $10^{1}$ to $10^{2}$, which is reasonable to evaluate the MSE but still has a gap between MSE and bound. As a result, when the compatibility constant is sufficiently large, the MSE would be small so that the lasso might perform well. Even if the compatibility constant is small, there exists situations where the MSE is sufficiently small compared with its bound.
@@ 4.2 Computation of upper bound via MIQP
The ratio remains almost the same across different time limits for a fixed $K$, indicating that the warm start provides a good initial value. In particular, when $K$ is large, the ratio becomes quite small. We also observe that the effect of $K$ is much larger than that of the time limit. This behavior suggests that most of the improvement can be almost achieved by the warm start based on $K$ random signs, and MIQP has only a limited search space left to explore. Therefore, increasing the time limit gives almost no additional benefit, while increasing $K$ significantly enhances the quality of the warm start. Thus, it is preferable to choose a large $K$, while the time limit does not necessarily have to be large; 60 seconds appears to be sufficient when $K$ is large.
Next, we consider a larger problem with $s=50$, where the QP with all sign combinations cannot be performed in practice due to the high computational cost of $2^{50-1}$ patterns. We compute an upper bound of $\phi$ using MIQP described above. Based on Figure 4, we set $K=20$ and a time limit of $T=60$ seconds. The other settings are the same as in Figure 1. The results are shown in Figure 5. The red dashed line indicates the compatibility constant for the population covariance matrix, $1-\rho$, while the black dashed line indicates $\phi=0$.
Figure 5: Compatibility constant $\phi$ with $s=50$ with for varying $(n,p,\rho)$. The red dashed line indicates the compatibility constant for the population covariance matrix, $\sqrt{1-\rho}$, while the black dashed line indicates $\phi=0$.
The results show similar tendency to Figure 1, but yields much lower $\phi$ for small sample sizes. In particular, when $p=5000$ and $n=1000$ (high-dimensional and relatively large sample sizes), the upper bound of compatibility constant is 0, indicating that the compatibility constant is also 0. Therefore, even when the number of observations is large, the compatibility constant can be zero or close to zero when $s$ is relatively large. This result shows a clear finite-sample gap compared with asymptotic intuition. More surprisingly, even in the uncorrelated case ($\rho=0$), the compatibility condition fails simply due to the relationship among $(n,p,s)$. Therefore, the finite-sample behavior is strongly dominated by these three factors rather than true correlation among predictors. The MIQP with time limit cannot always be exact, but is helpful here because it can detect that $\phi$ becomes zero or close to zero in a challenging settings. This observation cannot be obtained by QP alone due to its high computational cost.
@@ 5 Real data analysis
To compute the compatibility constant, we need the zero/nonzero pattern of the true coefficients, which cannot be provided in practice. Therefore, we provide a heuristic method to estimate the zero/nonzero pattern of the true coefficients. First, we divide the dataset into training and test datasets. The number of observations for the training and test data is $n_{\rm train}=1009$ and $n_{\rm test}=1000$, respectively. Then, we fit the lasso and estimate the set $S$ as follows:
@@
    We first select the regularization parameter $\lambda$ by the 10-fold cross validation. The cross-validation is known to be inconsistent in model selection and tends to select overly complex models (shao1993linear). Indeed, the number of nonzero coefficients selected by the cross-validation is $77$, which is relatively large.
@@
    The fitted value from cross-validation is expected to be sufficiently small because cross-validation is an unbiased estimator of the prediction error. Therefore, the unbiased estimator of $\sigma^{2}$, $\hat{\sigma}^{2}=\frac{1}{n_{\mathrm{train}}-s_{\mathrm{CV}}-1}\bigl\|\bm{y}_{\mathrm{train}}-X_{\mathrm{train}}\hat{\bm{\beta}}^{(\mathrm{CV})}\bigr\|_{2}^{2},$ would be a good estimator. Numerically, it is $\hat{\sigma}^{2}=1.268\times 10^{-4}.$
@@
The estimate of the active set, $\hat{S}$, can be used to compute the compatibility constant with the test dataset. To investigate the behavior of the compatibility constant with various sample sizes, we use the first $n=100k$ ($k=1,2,\dots,10$) observations of the test block. Denote the resulting subsample by $(X_{n},y_{n})$. For each $n$, we compute the empirical compatibility constant of $X_{n}$ with respect to the fixed set $\hat{S}$, say $\phi_{n}=\phi\bigl(\hat{S};X_{n}\bigr)$, using OSQP.
We also estimate the MSE and its bound computed with the $\phi_{n}$ to obtain the results similar to the Figures 13 in the simulation study. For a given $n$, we define the regularization parameter as $\lambda_{n}=2\hat{\sigma}\sqrt{\frac{2}{n}\Bigl(1+\log\frac{p}{\delta}\Bigr)},$ with $\delta=0.1$, which is the analog of (14) with $n$ in place of $n_{\mathrm{train}}$. We then compute the lasso estimate, $\hat{\bm{\beta}}_{n}$, with the above regularization parameter. The MSE in (12) is then estimated as
$\widehat{\mathrm{MSE}}=\frac{1}{n}\bigl\|X_{n}(\hat{\bm{\beta}}_{n}-\hat{\bm{\beta}}^{\rm train})\bigr\|_{2}^{2},\quad\mathrm{Bound}_{n}:=\frac{9\,|\hat{S}|\,\lambda_{n}^{2}}{\phi_{n}^{2}}.$
@@
Figure 7 shows the estimate of compatibility constant $\phi_{n}$, the MSE and its error bound scaled by $n/\log p$ as described in (13) in the simulation study, and their ratio, $\mathrm{Ratio}_{n}={\mathrm{Bound}_{n}}/\widehat{\mathrm{MSE}}$.
Figure 7: The estimate of compatibility constant $\phi_{n}$ (left panel), the MSE and its error bound scaled by $n/\log p$ as described in (13) in the simulation study (middle panel), and their ratio, $\mathrm{Ratio}_{n}={\mathrm{Bound}_{n}}/\widehat{\mathrm{MSE}}$ (right panel).
The compatibility constant is a monotone non-decreasing function of $n$, which implies a large sample size is required to achieve a good error bound. For $n=100$ and $n=200$, $\phi_{n}=0$ and then the compatibility condition may not hold. Meanwhile, for large $n$, the compatibility constant can be greater than $0.3$, but it is still not very large. Considering that there exist a few pairs of large correlations in predictors, as shown in Figure 6, the moderate value of $\phi_{n}$ seems reasonable.
The scaled $\widehat{\mathrm{MSE}}$ is almost constant as a function of $n$, while the bound is a monotone decreasing function, which aligns with the simulation results. This monotonicity is likely due to the fact that the compatibility constant is non-decreasing in $n$. The error bounds for $n=100$ and $200$ cannot be obtained because the compatibility condition does not hold. The ratio is almost a monotone decreasing function, which also align with the results from simulation study. The ratio is not small but not too large compared to the results from the simulation study (see Figure 3). As a result, the overall results almost align with the simulation study and may provide relatively good accuracy when a sufficient number of observations is available.

### v2

@@ 3.2 Mixed-Integer Quadratic Programming
The problem (8) can be easily solved when the number of elements in active set, $s$, is sufficiently small. However, since the number of sign patterns in (8) is $2^{s-1}$ and grows exponentially, the computation of (8) becomes infeasible when $s$ is large.
To avoid enumerating all $2^{s-1}$ sign patterns explicitly, we formulate the problem as a mixed-integer quadratic program (MIQP, Bertsimas and Tsitsiklis, 1997; Bertsimas and Weismantel, 2005). For each $j\in S$, we introduce nonnegative variables $v_{j}^{+},v_{j}^{-}\geq 0$ and a binary variable $b_{j}\in\{0,1\}$ to express $v_{j}$ as follows:
$v_{j}=v_{j}^{+}-v_{j}^{-},\quad v_{j}^{+}\leq Mb_{j},\quad v_{j}^{-}\leq M(1-b_{j}),$
where $M>0$ is a sufficiently large constant. This formulation is referred to as Big-$M$ formulation (Nemhauser and Wolsey, 1988). With this formulation, the constraint $\|\bm{v}_{S}\|_{1}=1$ can be written as a linear equality
@@
Alternatively, we may replace the Big-$M$ constraints $v_{j}^{+}\leq Mb_{j}$ and $v_{j}^{-}\leq M(1-b_{j})$ in (9) with the SOS1 (special ordered set of type 1) constraint on $(v_{j}^{+},v_{j}^{-})$, which enforces that at most one of them is nonzero (Beale and Tomlin, 1970; Williams, 2013).
@@ 4.1 Evaluation of compatibility constant via QP
For each iteration of the Monte Carlo simulation, we generate a dataset and compute the compatibility constant $\phi$ by solving the QP with the OSQP solver using the osqp package in R. The lasso estimator is implemented with glmnet. We obtain the following mean squared prediction error (MSPE) and its upper bound:
$\mathrm{MSPE}=\frac{1}{n}\|\bm{X}(\hat{\bm{\beta}}-\bm{\beta})\|_{2}^{2},\qquad\mathrm{bound}=9\lambda^{2}s/\phi^{2}.$ (10)
@@
Figure 1 presents the compatibility constant $\phi$ for different $n$, $p$, and $\rho$ under $R=10$ runs. The results show that the compatibility constant becomes larger as the number of observations increases. This behavior can be explained by the fact that the rank of the Gram matrix $\hat{\Sigma}=\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}/n$ is less than or equal to $\min(n,p)$. When $n$ is small, the Gram matrix $\hat{\Sigma}$ has low rank and a large null space, allowing vectors satisfying the condition $\mathcal{C}:=\{\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}\}$ to align with directions where the quadratic form $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ becomes nearly zero. As $n$ increases, the rank of $\hat{\Sigma}$ increases and its null space shrinks, and as a result, it becomes harder to find such directions within $\mathcal{C}$.
@@
We also observe that as $p$ increases, the compatibility constant decreases. This is probably because increasing $p$ enlarges the feasible set $\mathcal{C}$, which provides more degrees of freedom and makes it easier to find directions in which the quadratic form $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ is small. In particular, when $p$ is large and $n$ is small, the compatibility constant sometimes becomes $\phi=0$, which implies the compatibility condition does not hold. Therefore, for high-dimensional data, several hundred observations would be necessary to get a reasonable compatibility constant. The compatibility constant also highly depends on the value of $\rho$; as the value of $\rho$ increases, the value of $\phi$ decreases. This can be explained by noting that the compatibility constant behaves similarly to the square root of the smallest eigenvalue of the covariance matrix, $\sqrt{1-\rho}$, which decreases as $\rho$ increases.
@@
The red dashed line indicates the theoretical upper bound for the compatibility constant derived from the population covariance matrix, $\Sigma$; when $s$ is even, $\phi^{2}=1-\rho$, and when $s$ is odd, $\phi^{2}\leq(1-\rho)\left(1+\frac{1}{s(p-s)}\right)$. For the derivation of these upper bounds, please refer to Appendix A. We observe that when $p$ is small, the compatibility constant converges to the upper bound as $n$ becomes large. However, when $p$ is large, the speed of convergence to the bound is slow. In other words, even under moderate correlation structures, the sample compatibility constant can remain far from the population-based bound because of finite-sample limitations rather than the population correlation itself.
Figure 2: Scaled MSPE and upper bound, ${\rm MSPE}_{\rm scaled}$ and ${\rm bound}_{\rm scaled}$, for varying $(n,p,\rho)$.
Figure 2 shows the scaled MSPE and its upper bound as below
${\rm MSPE}_{\rm scaled}=\frac{n}{\log p}{\rm MSPE},\quad{\rm bound}_{\rm scaled}=\frac{n}{\log p}{\rm bound},$ (11)
@@
where the MSPE and its bound are defined in (10). Here, the “scaled” means multiplication of $n/\log p$ so that the right-hand side of inequality (4), ${\rm MSPE}\lesssim C\log p/n$, has a quantity of order one.
@@
The result shows that the scaled MSPE is relatively stable regardless of the value of $n$ and $p$, except for small $n$. On the other hand, the upper bound is unstable when $n$ is small and $p$ is large. This is because the compatibility constant can be small for small $n$ as discussed in Figure 1. In particular, with large $p$ small $n$, the upper bound cannot be obtained because of $\phi=0$. Therefore, the upper bound made by the small $n$ can be much larger than that with large $n$.
@@
Figure 3: Ratio between the MSPE and the upper bound, $\text{MSPE}/\text{Bound}$, for varying $(n,p,\rho)$.
@@
Figure 3 shows the ratio between the MSPE and its upper bound. The ratio becomes extremely small for high-dimensional settings and small sample sizes. In particular, the ratio can be $10^{-6}$ to $10^{-5}$, which implies that the upper bound does not provide a meaningful evaluation of the MSPE. Therefore, the compatibility constant cannot be useful unless $n$ is relatively large. For large sample sizes, the ratio is $10^{-2}$ to $10^{-1}$, which is reasonable to evaluate the MSPE but still has a gap between MSPE and bound. As a result, when the compatibility constant is sufficiently large, the MSPE would be small so that the lasso might perform well. Even if the compatibility constant is small, there exist situations where the MSPE is sufficiently small compared with its bound.
4.2 Evaluation of the prediction error bound at each step in the proof of the oracle inequality
In the previous simulation study, we observe that the gap between the MSPE and its bound is generally large, especially when $p$ is large. The gap comes from many inequalities obtained from each step of the proof in Proposition 4. The inequality becomes looser as the steps of the proof increase. Here, we evaluate the gap for each step of the proof, and investigate the key inequality that loosens the gap significantly.
Table 1: Bound of MSPE for each step of the proof. Here, $\bm{\Delta}:=\hat{\bm{\beta}}-\bm{\beta}$.
Step Inequality Bound
B Basic inequality $\mathrm{MSPE}\leq\frac{2}{n}\bm{\varepsilon}^{\top}X\bm{\Delta}+2\lambda\bigl(\|\bm{\beta}\|_{1}-\|\hat{\bm{\beta}}\|_{1}\bigr)$
Du Dual norm $\mathrm{MSPE}\leq 2\left\|\frac{X^{\top}\bm{\varepsilon}}{n}\right\|_{\infty}\|\bm{\Delta}\|_{1}+2\lambda\bigl(\|\bm{\beta}\|_{1}-\|\hat{\bm{\beta}}\|_{1}\bigr)$
M Max bounded by $\lambda/2$ $\mathrm{MSPE}\leq\lambda\|\bm{\Delta}\|_{1}+2\lambda\bigl(\|\bm{\beta}\|_{1}-\|\hat{\bm{\beta}}\|_{1}\bigr)$
T Triangle $\mathrm{MSPE}\leq 3\lambda\|\bm{\Delta}_{S}\|_{1}-\lambda\|\bm{\Delta}_{S^{c}}\|_{1}$
Dr Drop negative term $\mathrm{MSPE}\leq 3\lambda\|\bm{\Delta}_{S}\|_{1}$
C Compatibility $\mathrm{MSPE}\leq 3\lambda\sqrt{\frac{s}{\phi^{2}}}\sqrt{\mathrm{MSPE}}$
F Final $\mathrm{MSPE}\leq\frac{9\lambda^{2}s}{\phi^{2}}$
The proof of the oracle inequality in (3) mainly consists of 7 steps. For detail, please refer to Appendix B. Table 1 shows the bounds made by each step of the inequality. First, we employ a basic inequality (referred to as “B”) constructed by the minimization of the empirical loss. Next, the inequality $|\bm{a}^{T}\bm{b}|\leq\|\bm{a}\|_{\infty}\|\bm{b}\|_{1}$ for given two vectors, $\bm{a}$ and $\bm{b}$, is used to evaluate the inner product of error variables and the prediction values (referred to as “Du” because it involves the dual inequality). The bound, $\left\|X^{\top}\bm{\varepsilon}\right\|_{\infty}$, made by the dual inequality is bounded by $n\lambda/2$ with probability at least $1-\delta$ (referred to as “M”, indicating Maximum bound). We then employ the triangle inequality (referred to as “T”) and drop a negative term (referred to as “Dr”). The compatibility condition (referred to as “C”) is used to evaluate the value of $\|\bm{\beta}_{S}-\hat{\bm{\beta}}_{S}\|_{1}$. Finally, the oracle inequality is derived by squaring both sides of the inequality derived from the compatibility condition (referred to as “F”).
Figure 4: Ratio between the actual MSPE and its bound for each step of the proof of inequalities. The seven steps are denoted by B (basic inequality), Du (dual norm), M (max inequality), T (triangle inequality), Dr (dropping the negative term), C (compatibility inequality), and F (final bound).
Figure 4 shows the ratio between the MSPE and its bound for each step of the inequalities. For clarity, we consider sample sizes $n=500,2000$, numbers of variables $p=50,500,5000$, and correlation parameters $\rho=0$, $0.4$, and $0.8$. When evaluating the ratio, we first check the empirical frequency of the event, $\left\|X^{\top}\bm{\varepsilon}/n\right\|_{\infty}\leq\lambda/2.$ Theoretically, this event occurs with probability at least $1-\delta=0.9$, and we observe that it holds in all trials for almost all settings, except for $(n,p)=(2000,50)$ and $(2000,5000)$ with $\rho=0.4$, where the empirical frequency of this event is 0.9 over $R=10$ runs. The plots in Figure 4 are shown only for trials in which this event holds.
We obtain the following observations from Figure 4:
  • The bounds of T (triangle) and Dr (drop the negative term) do not decrease much for all settings. The bound of M (max bounded by $\lambda/2$) theoretically decreases with probability at least $1-\delta$, and it actually decreases in many cases but does not decrease much, except for the large $\rho$, large $p$ case.
  • B, Du, C, and F can result in a substantial decrease in the ratio. Thus, the looseness of the final bound is attributed to some of B, Du, C, and F.
  • For any cases, the ratio at B is often around one half. Indeed, the ratio is exactly the half when OLS (ordinary least squares), $\hat{\bm{\beta}}^{\rm OLS}$, is used. As $\bm{y}-X\hat{\bm{\beta}}^{\rm OLS}$ and $X(\hat{\bm{\beta}}^{\rm OLS}-\bm{\beta})$ are orthogonal, we have $\|\bm{y}-X\bm{\beta}\|^{2}=\|\bm{y}-X\hat{\bm{\beta}}^{\rm OLS}\|^{2}+\|X(\hat{\bm{\beta}}^{\rm OLS}-\bm{\beta})\|^{2}$. Consequently, ${\rm MSPE}=\bm{\varepsilon}^{\top}X(\hat{\bm{\beta}}^{\rm OLS}-\bm{\beta})/n$. The half ratio suggests that the lasso estimator behaves similarly to OLS, in the sense that $\bm{y}-X\hat{\bm{\beta}}$ and $X(\hat{\bm{\beta}}-\bm{\beta})$ are nearly orthogonal.
  • When $p$ and $\rho$ are small and $n$ is large, the bound made by the compatibility condition (C) is reasonably tight. However, outside such settings, the drop with C becomes much worse.
  • In all cases, the ratio drops substantially at F in the end. The final step removes the square root by squaring both sides, which substantially enlarges the bound.
@@ 4.3 Computation of upper bound via MIQP
The ratio does not change substantially across different time limits for a fixed $K$, indicating that the warm start provides a reasonable initial value. We also observe that the effect of $K$ is much larger than that of the time limit; as $K$ increases, the ratio decreases substantially. These findings suggest that most of the improvement is achieved by the warm start based on $K$ random signs, and MIQP has only a limited search space left to explore. Therefore, increasing the time limit gives almost no additional benefit, while increasing $K$ significantly enhances the quality of the warm start. It would be preferable to choose a large $K$, while the time limit does not necessarily have to be large; 60 seconds appears to be sufficient when $K$ is large.
Next, we consider a larger problem with $s=50$, where the QP with all sign combinations cannot be performed in practice due to the high computational cost of $2^{50-1}$ patterns. We compute an upper bound of $\phi$ using MIQP described above. Based on Figure 5, we set $K=20$ and a time limit of $T=60$ seconds. The other settings are the same as in Figure 1. The results are shown in Figure 6. The red dashed line indicates the compatibility constant for the population covariance matrix, $\sqrt{1-\rho}$, while the black dashed line indicates $\phi=0$.
Figure 6: Compatibility constant $\phi$ with $s=50$ for varying $(n,p,\rho)$. The red dashed line indicates the compatibility constant for the population covariance matrix, $\sqrt{1-\rho}$, while the black dashed line indicates $\phi=0$.
The results show a similar tendency to Figure 1, but yield much lower $\phi$ for small sample sizes. In particular, when $p=5000$ and $n=1000$ (high-dimensional and relatively large sample sizes), the upper bound of compatibility constant is 0, indicating that the compatibility constant is also 0. Therefore, even when $n$ is relatively large, the compatibility constant can be zero or close to zero. This result shows a clear finite-sample gap compared with asymptotic intuition. More surprisingly, even in the uncorrelated case ($\rho=0$), the compatibility condition does not hold simply due to the relationship among $(n,p,s)$. Therefore, the finite-sample behavior is strongly dominated by these three factors rather than true correlation among predictors. The MIQP with time limit cannot always be exact, but is helpful here because it can detect that $\phi$ becomes zero or close to zero in challenging settings.
@@ 5 Real data analysis
To compute the compatibility constant, we need the support of the true coefficients, which cannot be provided in practice. Therefore, we provide a heuristic method to estimate the support of the true coefficients. First, we divide the dataset into training and test datasets. The number of observations for the training and test data is $n_{\rm train}=1009$ and $n_{\rm test}=1000$, respectively. Then, we fit the lasso and estimate the set $S$ as follows:
@@
    We first select the regularization parameter $\lambda$ by the 10-fold cross validation. The cross-validation is known to be inconsistent in model selection and tends to select overly complex models (Shao, 1993). Indeed, the number of nonzero coefficients selected by the cross-validation is $77$, which is relatively large.
@@
    We obtain an estimator of $\sigma^{2}$ defined by $\hat{\sigma}^{2}=\frac{1}{n_{\mathrm{train}}-s_{\mathrm{CV}}-1}\bigl\|\bm{y}_{\mathrm{train}}-X_{\mathrm{train}}\hat{\bm{\beta}}^{(\mathrm{CV})}\bigr\|_{2}^{2}$. Numerically, $\hat{\sigma}^{2}=1.268\times 10^{-4}.$
@@
The estimate of the active set, $\hat{S}$, can be used to compute the compatibility constant with the test dataset. To investigate the behavior of the compatibility constant with various sample sizes, we use the first $n=100k$ ($k=1,2,\dots,10$) observations of the test block. Denote the resulting subsample by $(X_{n},y_{n})$. For each $n$, we compute the empirical compatibility constant of $X_{n}$ with respect to the fixed set $\hat{S}$, say $\phi_{n}=\phi\bigl(\hat{S};X_{n}\bigr)$, using QP.
We also estimate the MSPE and its bound computed with the $\phi_{n}$ to obtain the results similar to the Figures 13 in the simulation study. For a given $n$, we define the regularization parameter as $\lambda_{n}=2\hat{\sigma}\sqrt{\frac{2}{n}\Bigl(1+\log\frac{p}{\delta}\Bigr)},$ with $\delta=0.1$, which is the analog of (12) with $n$ in place of $n_{\mathrm{train}}$. We then compute the lasso estimate, $\hat{\bm{\beta}}_{n}$, with the above regularization parameter. The MSPE in (10) is then estimated as
$\widehat{\mathrm{MSPE}}=\frac{1}{n}\bigl\|X_{n}(\hat{\bm{\beta}}_{n}-\hat{\bm{\beta}}^{\rm train})\bigr\|_{2}^{2},\quad\mathrm{Bound}_{n}:=\frac{9\,|\hat{S}|\,\lambda_{n}^{2}}{\phi_{n}^{2}}.$
@@
Figure 8 shows the estimate of compatibility constant $\phi_{n}$, the MSPE and its error bound scaled by $n/\log p$ as described in (11) in the simulation study, and their ratio, $\mathrm{Ratio}_{n}=\widehat{\mathrm{MSPE}}/{\mathrm{Bound}_{n}}$.
Figure 8: The estimate of compatibility constant $\phi_{n}$ (left panel), the MSPE and its error bound scaled by $n/\log p$ as described in (11) in the simulation study (middle panel), and their ratio, $\mathrm{Ratio}_{n}=\widehat{\mathrm{MSPE}}/{\mathrm{Bound}_{n}}$ (right panel).
The estimated compatibility constant is a monotone non-decreasing function of $n$, which implies a large sample size is required to achieve a good error bound. For $n=100$ and $n=200$, $\phi_{n}=0$ and then the compatibility condition may not hold. Meanwhile, for large $n$, the compatibility constant can be greater than $0.3$, but it is still not very large. Considering that there exist a few pairs of large correlations in predictors, as shown in Figure 7, the moderate value of $\phi_{n}$ seems reasonable.
The scaled $\widehat{\mathrm{MSPE}}$ is almost constant as a function of $n$, while the bound is a monotone decreasing function, which aligns with the simulation results. This monotonicity is likely due to the fact that the compatibility constant is non-decreasing in $n$. The error bounds for $n=100$ and $200$ cannot be obtained because the compatibility condition does not hold. The ratio is a monotone increasing function, which also aligns with the results from simulation study. The ratio is not small but not too large compared to the results from the simulation study (see Figure 3). As a result, the overall results almost align with the simulation study and may provide relatively good upper bound when a sufficient number of observations is available.